OpenStudy (geometry_hater):

Area of the square P is 36 square inches and the perimeter of the rectangle R is 144 inches. If a width of the rectangle is two times a length of the square, which is the length of a length of the rectangle? Answer 70 60 45 30

OpenStudy (geometry_hater):

i dont know how to do this

OpenStudy (anonymous):

60

OpenStudy (anonymous):

Yes 60

OpenStudy (geometry_hater):

but how?

OpenStudy (anonymous):

Side of square = 6 Width of rectangle = 2*6=12 2(l+12)=144 2l+24=144 2l=120 l=60

OpenStudy (anonymous):

p= 36 right? Square root = 6 6= length of square, 12 is length of triangle, because it's two times longger 12 times two subtracted from 144, then divided by two is sixty

OpenStudy (anonymous):

*rectangle